OpenStudy (anonymous):

Factor by grouping, if possible, and check. 5x3 + 4x2 − 10x − 8

OpenStudy (anonymous):

okay am a bit lost this is new for me so how did you get the answer that you did

OpenStudy (anonymous):

\[5x^3 +4x^2 - 10x -8 = x^2(5x+4) -2(5x+4) = (x^2-2)(5x+4)\]

OpenStudy (anonymous):

ok